# ChangeMamba: Remote Sensing Change Detection with Spatio-Temporal State Space Model

### v1

@@ II-C State Space Model
The SSM concept first popped up with the S4 model [28], offering a new way to handle contextual information globally, which caught attention because of its attractive property of scaling linearly in sequence length. Based on S4, Smith et al. [46] proposed a new S5 model by introducing MIMO SSM and efficient parallel scan into the S4 model. After that, the H3 model [47] further advanced these concepts, achieving performance on par with Transformers in language modeling tasks. Recently, Gu et al [29] proposed a data-dependent SSM layer and built a generic language model backbone called Mamba, which outperforms Transformers at at different scales on large-scale real data and scales linearly in sequence length. Very recently, VMamba [30] and Vision Mamba [31] have extended Mamba architecture to 2D image data, showing superior performance on many computer vision tasks. Inspired by this progress, some pioneering work has introduced the Mamba architecture to the field of remote sensing image processing [48, 49]. Their studies show that the Mamba architecture can yield comparable performance to the advanced CNN and Transfomer architectures on scene classification [49] and pansharpening [48] tasks. However, these works still focus on either low-level tasks or classification tasks, and all of these tasks are single-temporal tasks [50]. The potential of the Mamba architecture for multi-temporal remote sensing image-related scenarios and intensive prediction tasks remains to be explored.
@@ III-A Preliminaries
The S4 and Mamba models are discrete counterparts of the continuous system as continuous systems face significant challenges when integrated into deep learning algorithms. They contain a time scale parameter which is used to convert the continuous parameters and into their discrete counterparts and . A prevalent way to achieve this transformation is to use a zero-order hold (ZOH) approach:
@@
Finally, the output of with sequence length can be calculated directly by the following global convolution operation
@@ III-B1 Binary Change Detection
BCD is the basic and the most intensively studied task in the CD field. It focuses on “where” change occurs. According to the class of interest, BCD can also be categorized into class-agnostic CD focusing on general land-cover change information, and single-class object CD, e.g., building CD. Given a training set represented as , where is the -th multi-temporal image pair acquired in and , respectively, and is the corresponding label, the goal of BCD is to train a change detector on that can predict change maps reflecting accurate “change / non-change” binary information on new sets.
@@ III-B3 Building Damage Assessment
BDA is a special “one-to-many” SCD task [3]. BDA requires recognizing not only the where of the change (damage) occurs but also post-event states of the object (building) in the condition that the pre-event state of the object is unique. The training set in BDA task can be denoted as , where is the label of the object of interest (building) at , and is the post-event state (damage level) of the object of interest at .
@@ III-C Network Architecture
For these three CD subtasks, we design corresponding network frameworks based on the Mamba architecture, called MambaBCD, MambaSCD, and MamabBDA, respectively. Fig. 1 shows the network framework of these three architectures. Among them, the encoders of the three networks are weight-sharing siamese networks based on the Visual State Space Model (VMamba) architecture [30]. VMamba can adequately extract the robust and representative features of the input images benefiting from the Mamba architecture and an efficient 2D cross-scan mechanism (as shown in Fig. 3). Its specific structure will be introduced in Section III-D.
@@ III-C1 MambaBCD
MambaBCD is the architecture designed for the BCD task. First, the siamese encoder network extracts multi-level features from the input image as and . Next, these multi-level features are fed into a tailored change decoder . Based on the Mamba architecture, the change decoder can fully learn the spatio-temporal relationship from multi-level features through three different mechanisms, and gradually obtains an accurate BCD result, formulated as . The binary change map can be obtained as .
@@ III-C2 MambaSCD
MambaSCD is the architecture designed for SCD tasks. As shown in Fig. 1-(b), MambaSCD adds two semantic decoders for land cover mapping tasks based on MambaBCD, denoted as and . In addition to be treated as input as to learn spatio-temporal relationships to predict BCD results, the multi-level features extracted by the encoder are also fed into and to predict the land-cover map of the corresponding temporal image, formulated as and . After obtaining the land-cover maps and and the binary change map , the semantic change information of can be obtained by performing the mask operation on and using .
@@ III-D Encoder Based on Visual State Space Model
S6 model in Mamba causally processes the input data, and thus can only capture information within the scanned part of the data. This attribute naturally aligns with natural language data, which but poses significant challenges when adapting to non-causal data such as image. Directly expanding image data along a certain dimension of space and inputting it into the S6 model will result in spatial context information not being adequately modeled. Recently, visual Mamba solved this problem well by proposing a 2D Cross-Scan mechanism. As shown in Fig. 3, before inputting the tokens into the S6 model, the cross-scan mechanism rearranges the tokens in spatial dimensions, i.e., top-left to bottom-right, bottom-right to top-left, top-right to bottom-left, and bottom-left to top-right. Finally, the resulting features are then merged. In this way, any pixel can get spatial context information from different directions. Moreover, the computational complexity of the Mamba model under the cross-scan mechanism is still compared to the self-attention technique in Transformer.
Based on the Visual Mamba architecture, the specific structure of the encoder network in our proposed three architectures is shown in Fig. 2. There are four stages, each of which first downsamples the input data, then fully models the spatial contextual information using a number of visual state space (VSS) blocks, and then outputs the features for that stage, i.e., . The structure of VSS block is also depicted in Fig. 2. The input first passes through a linear embedding layer, and the output splits into two information flows. One flow passes through a 3 3 depth-wise convolution (DWConv) layer, followed by a Silu activation function [52] before entering the core SS2D module (i.e., the integration of S6 with the cross-scan mechanism). The output of the SS2D module passes through a layer normalization (LN) layer, and is then summed up with the outputs of the other streams that have been activated by Silu. This combination produces the final output of the VSS block. Finally, the features from the four stages are then used in the subsequent decoders responsible for specific tasks.
@@ III-E1 Spatio-Temporal Relationship Modelling Mechanism
Although the employed encoder can extract robust and representative features, learning spatio-temporal relationships of multi-temporal images is also significant to CD tasks [36, 18]. To this end, we propose three mechanisms for modeling spatio-temporal relationships that can be aligned with the attribute of the S6 model capable of adequately modeling the global contextual information of long sequential data. Fig. 4 illustrates the three modeling mechanisms. They are spatio-temporal sequential modelling, spatio-temporal cross modeling, and spatial-temporal parallel modeling. Sequential modelling is unfolding the tokens of two temporal phases of data and then arranging them in temporal order as . Cross modelling mechanism, on the other hand, is modeling by cross-ordering the tokens of the two temporal phases as . Finally, parallel modeling is done by concatenating the tokens of the two temporal phases in the channel dimension and then performing joint modeling, i.e., . Through these three mechanisms and Mamba architecture, the spatio-temporal relationships intrinsic in the multi-temporal features will be fully explored, helping the decoder to obtain accurate change detection results.
@@ III-E2 Change Decoder
Based on the proposed three spatio-temporal learning mechanisms, the specific structure of the change decoder is then shown in Fig. 5. It fully learns the spatio-temporal relationships from the extracted multi-temporal features and in four stages and obtains accurate binary change maps. At the beginning of each stage, the spatio-temporal relationship of multi-temporal features is first modeled using an STSS block. In the STSS block, a spatio-temporal tokens generator module will rearrange the input multi-temporal features, which are then fed into three VSS blocks. Each block is responsible for learning one of the spatio-temporal relationships in Fig. 4. The output of the STSS block of the current stage is then integrated with the information from the feature map of the previous stage through a fusion module. After passing through an upsampling layer, the feature map is fed to the next stage.
@@ III-E3 Semantic Decoder
The specific structure of the semantic decoder is then shown in Fig. 6. It is mainly responsible for gradually recovering the class-agnostic or object-specific land-cover maps of the corresponding multi-level features extracted by the encoder. It also has four stages. At the beginning of each stage, the global spatial context information of input data is first modeled using a VSS block. Then the feature map is up-sampled and integrated with information about the lower-level feature map with higher resolution through a fusion module. In the fusion module, the number of channels of the low-level feature map is mapped to be consistent with the high-level feature map by a 11 convolutional layer. Then the high-level and low-level feature maps are summed. Finally, the resulting feature map is smoothed by a residual layer.
@@ III-F Loss Function
Since this paper focuses on exploring the potential of the Mamba architecture for CD, we have optimized networks using commonly used loss functions in the field. Some tricks for improving accuracy like focal loss [53], deep supervision [17], multi-scale training/testing [6], etc. are not adopted. Nevertheless, in the experimental section, we can find that the proposed architecture achieves very competitive and even SOTA performance on all three subtasks.
@@ III-F1 BCD
Since BCD can be regarded as a special semantic segmentation task [50], we directly optimize the network using cross-entropy loss as
@@ III-F3 BDA
BDA contains a building location task and a building damage classification task. The cross-entropy loss for each task can be expressed as follows
@@ IV-A1 SYSU-CD [15]
This dataset introduces a comprehensive collection of 20,000 pairs of 0.5-meter resolution aerial images from Hong Kong, spanning 2007 to 2014, to advance the field of CD. This dataset is distinguished by its focus on urban and coastal changes, featuring high-rise buildings and infrastructure developments, areas where CD poses significant challenges due to shadow and deviation effects. For effective deep learning (DL) application, the dataset is structured into training, validation, and test sets following a 6:2:2 ratio, with 20,000 patches of 256 × 256. Covering diverse change scenarios such as urban construction, suburban expansion, groundwork, vegetation changes, road expansion, and sea construction, the SYSU-CD dataset is positioned as a pivotal resource for CD research.
@@ IV-A2 WHU-CD [55]
The WHU-CD dataset, a segment of the larger WHU Building dataset, is tailored for the task of building change detection. It comprises two aerial datasets from Christchurch, New Zealand, captured in April 2012 and 2016, with a high image resolution of 0.3m. This dataset is particularly focused on detecting changes in large and sparse building structures. The 2012 dataset includes aerial imagery covering an area of 20.5 km², featuring 12,796 buildings, while the 2016 dataset shows an increase to 16,077 buildings within the same area, reflecting significant urban development over the four-year period. We cropped these large training (21,243 × 15,354) and testing (11,265 × 15,354) images into non-overlapping 256 × 256 patches for our experiments.
@@ IV-A3 LEVIR-CD+ [56]
This advanced version of the LEVIR-CD dataset introduces a comprehensive resource for BCD research. It comprises 985 pairs of very high-resolution Google Earth images at 0.5m/pixel resolution, each with dimensions of 1024 × 1024 pixels. Spanning a time interval of 5 to 14 years, these bi-temporal images document significant building construction changes. It also encompasses a wide array of building types, including villa residences, tall apartments, small garages, and large warehouses, with a focus on both the emergence of new buildings and the decline of existing structures. The dataset is meticulously annotated with binary labels by experts in remote sensing image interpretation. Each image pair undergoes a rigorous two-step quality assurance process involving initial annotation and subsequent verification to ensure the accuracy of change instances. The final dataset features 31,333 instances of building changes, making LEVIR-CD+ a valuable benchmark for evaluating CD methodologies.
@@ IV-A4 SECOND [38]
The SECOND dataset introduces a collection of 4,662 pairs of 512 × 512 aerial images at 0.5-3m/pixel resolution, annotated for SCD in cities such as Hangzhou, Chengdu, and Shanghai. It focuses on six primary land cover classes: non-vegetated ground surfaces, trees, low vegetation, water, buildings, and playgrounds to capture a wide range of geographical changes. Annotations are done by experts, ensuring high accuracy at the pixel level. The dataset uses land-cover map pairs and non-change masks to accurately represent change, facilitating the differentiation of changes from unchanged pixels within the same class. This methodological approach ensures its utility in addressing SCD challenges by providing a diverse and accurately labeled dataset for model training and evaluation.
@@
The xBD dataset, pivotal for advancing CD and BDA in disaster response, integrates pre- and post-disaster satellite imagery with over 850,000 building annotations across 45,362 km². It stands as the largest dataset of its kind, enriched by environmental data and detailed damage levels for a broad spectrum of disaster events. Developed in collaboration with disaster response agencies, xBD is engineered to refine emergency response strategies, including resource allocation and aid distribution. The dataset adheres to stringent criteria, featuring high-resolution imagery (below a 0.8m GSD), encapsulating multiple levels of damage, and covering a diverse range of disaster types and building structures. By including both damaged and undamaged sites, xBD is crafted to foster the creation of adaptive models that effectively navigate the complexities of various disaster scenarios.
| Dataset | Study site | Number of image pairs | Image size | Spatial resolution | Evaluation task |
| SYSU | Hong Kong | 20,000 | 256 × 256 | 0.5m | BCD (class-agnostic) |
| LEVIR-CD+ | Texas of the US | 985 | 1,024 × 1,024 | 0.5m | BCD (single-object) |
| WHU-CD | Christchurch, New Zealand | 2 | 21,243/11,265 × 15,354 | 0.3m | BCD (single-object) |
| xBD | 15 countries | 22,068 | 1,024 × 1,024 | <0.8m | BDA |

### v2

@@ II-C State Space Model
The SSM concept first popped up with the S4 model [36], offering a new way to handle contextual information globally, which caught attention because of its attractive property of scaling linearly in sequence length. Based on S4, Smith et al. [56] proposed a new S5 model by introducing MIMO SSM and efficient parallel scan into the S4 model. After that, the H3 model [57] further advanced these concepts, achieving performance on par with Transformers in language modeling tasks. Recently, Gu et al. [37] proposed a data-dependent SSM layer and built a generic language model backbone called Mamba, which outperforms Transformers at different scales on large-scale real data and scales linearly in sequence length. Very recently, VMamba [38] and Vision Mamba [39] have extended Mamba architecture to 2D image data, showing superior performance on many computer vision tasks. Inspired by this progress, some pioneering work has introduced the Mamba architecture to the field of remote sensing image processing [58, 59]. Their studies show that the Mamba architecture can yield performance comparable to the advanced CNN and Transformer architectures on scene classification [59] and pansharpening [58] tasks. However, these works still focus on either low-level tasks or classification tasks, and all of these tasks are single-temporal tasks [44]. The potential of the Mamba architecture for multi-temporal remote sensing image-related scenarios and intensive prediction tasks remains to be explored.
@@ III-A Preliminaries
The S4 model [36] is a discrete counterpart of the continuous system as continuous systems face significant challenges when integrated into deep learning algorithms. It contains a time scale parameter which is used to convert the continuous parameters and into their discrete counterparts and . A prevalent way to achieve this transformation is to use a zero-order hold (ZOH) approach:
@@
The output of with sequence length can be calculated directly by the following global convolution operation
@@
Mamba architecture [37] builds on S4 model by proposing a selection mechanism that allows the model to filter out irrelevant information and remember relevant information. In addition, a hardware-aware algorithm has been developed for recursive computation of the model, but will not be implemented in the extended state, optimizing the GPU memory layout. These characteristics allow the Mamba architecture not only to obtain promising results on languages and other data with long sequences but also to achieve fast training and inference with memory and computational overhead linearly proportional to sequence length.
@@ III-B1 Binary Change Detection
BCD is the basic and the most intensively studied task in the CD field. It focuses on “where” change occurs. According to the class of interest, BCD can also be categorized into category-agnostic CD focusing on general land-cover change information, and single-category CD, e.g., building CD. Given a training set represented as , where is the -th multi-temporal image pair acquired in and , respectively, and is the corresponding label, the goal of BCD is to train a change detector on that can predict change maps reflecting accurate “change / non-change” binary information on new sets.
@@ III-B3 Building Damage Assessment
BDA is a special “one-to-many” SCD task [3]. BDA requires recognizing not only the “where” of the change (damage) occurs but also post-event states of the object-of-interest (building) given the pre-event state of the object-of-interest. The training set in BDA task can be denoted as , where is the label of the object-of-interest (building) at , and is the post-event state (damage level) of the object of interest at .
@@ III-C Network Architecture
For these three CD subtasks, we design corresponding network frameworks based on the Mamba architecture, called MambaBCD, MambaSCD, and MamabBDA, respectively. Fig. 1 shows the network framework of these three architectures. Among them, the encoders of the three networks are weight-sharing siamese networks based on the Visual State Space Model (VMamba) architecture [38]. VMamba can adequately extract the robust and representative features of the input images benefiting from the Mamba architecture and an efficient 2D cross-scan mechanism (as shown in Fig. 3). All the three architectures have a change detector to learn spatio-temporal relationships from features extracted by the encoder. In addition, MambaSCD and MambaBDA additionally possess semantic decoders for other tasks. The specific structure of encoder and decoders in our architectures will be elaborated in Sections III-D and III-E. In this subsection, we focus on the input, output and internal information flow of these three network architectures.
@@ III-C1 MambaBCD
MambaBCD is the architecture designed for the BCD task. First, the siamese encoder network denoted as extracts multi-level features from the input multi-temporal images as and . Next, these multi-level features are fed into a tailored change decoder . Based on the Mamba architecture, the change decoder can fully learn the spatio-temporal relationship from multi-level features through three different mechanisms, and gradually obtains an accurate BCD result, formulated as . The binary change map can be obtained as .
@@ III-C2 MambaSCD
MambaSCD is the architecture designed for the SCD task. As shown in Fig. 1-(b), MambaSCD adds two additional semantic decoders for land cover mapping tasks based on MambaBCD, denoted as and . In addition to be treated as input as to learn spatio-temporal relationships to predict BCD results, the multi-level features extracted by the encoder are also fed into and to predict the land-cover map of the corresponding temporal image, formulated as and . After obtaining the land-cover maps and and the binary change map , the semantic change information of can be obtained by performing the mask operation on and using [45, 24, 6].
@@ III-D Encoder Based on Visual State Space Model
Mamba architecture causally processes the input data, and thus can only capture information within the scanned part of the data. This attribute naturally aligns with natural language data, which but poses significant challenges when adapting to non-causal data such as image [38]. Directly expanding image data along a certain dimension of space and inputting it into the S6 model will result in spatial contextual information not being adequately modeled. Recently, Visual Mamba solved this problem well by proposing a 2D Cross-Scan mechanism. As shown in Fig. 3, before inputting the tokens into the S6 model, the cross-scan mechanism rearranges the tokens in spatial dimensions, i.e., top-left to bottom-right, bottom-right to top-left, top-right to bottom-left, and bottom-left to top-right. Finally, the resulting features are then merged. In this way, any pixel can get spatial context information from different directions. Moreover, the computational complexity of the Visual Mamba model under the cross-scan mechanism is still compared to the self-attention operation in Transformer.
Based on the Visual Mamba architecture, the specific structure of the encoder network in our proposed three architectures is shown in Fig. 2. There are four stages, each of which first downsamples the input data, then fully models the spatial contextual information using a number of visual state space (VSS) blocks, and then outputs the features for that stage, i.e., . The structure of VSS block is also depicted in Fig. 2. The input first passes through a linear embedding layer, and the output splits into two information flows. One flow passes through a 3 3 depth-wise convolution (DWConv) layer, followed by a Silu activation function [61] before entering the core SS2D module (i.e., the integration of S6 model with the cross-scan mechanism). The output of the SS2D module passes through a layer normalization (LN) layer, and is then summed up with the outputs of the other streams that have been activated by Silu. This combination produces the final output of the VSS block. Finally, the features from the four stages are then used in the subsequent decoders responsible for specific tasks.
@@ III-E1 Spatio-Temporal Relationship Modelling Mechanism
Although the employed encoder can extract robust and representative features, learning spatio-temporal relationships of multi-temporal images is also significant to CD tasks [46, 24]. To this end, we propose three mechanisms for modeling spatio-temporal relationships that can be aligned with the attribute of the S6 model capable of adequately modeling the global contextual information of long sequential data. Fig. 5 illustrates the three modeling mechanisms. They are spatio-temporal sequential modelling, spatio-temporal cross modeling, and spatial-temporal parallel modeling. Sequential modelling is unfolding the tokens of two temporal phases of data and then arranging them in temporal order as . Cross modelling mechanism, on the other hand, is modeling by cross-ordering the tokens of the two temporal phases as . Finally, parallel modeling is done by concatenating the tokens of the two temporal phases in the channel dimension and then performing joint modeling, i.e., . Through these three mechanisms and Mamba architecture, the spatio-temporal relationships intrinsic in the multi-temporal features will be fully explored, helping the change decoder to obtain accurate change detection results.
@@ III-E2 Change Decoder
Based on the proposed three spatio-temporal learning mechanisms, the specific structure of the change decoder is then shown in Fig. 4. It fully learns the spatio-temporal relationships from the extracted multi-temporal features and in four stages and obtains accurate binary change maps. At the beginning of each stage, the spatio-temporal relationship of multi-temporal features is first modeled using an spatio-temporal state space (STSS) block. In the STSS block, a spatio-temporal tokens generator module will rearrange the input multi-temporal features, which are then fed into three VSS blocks. Each block is responsible for learning one of the spatio-temporal relationships in Fig. 5. The output of the STSS block of the current stage is then integrated with the information from the feature map of the previous stage through a fusion module. After passing through an upsampling layer, the feature map is fed to the next stage.
@@ III-E3 Semantic Decoder
The specific structure of the semantic decoder is then shown in Fig. 6. It is mainly responsible for gradually recovering the land-cover maps from the corresponding multi-level features extracted by the encoder. It also has four stages. At the beginning of each stage, the global spatial context information of input data is first modeled using a VSS block. Then the feature map is up-sampled and integrated with information about the lower-level feature map with higher resolution through a fusion module. In the fusion module, the number of channels of the low-level feature map is mapped to be consistent with the high-level feature map by a 11 convolutional layer. Then the high-level and low-level feature maps are summed. Finally, the resulting feature map is smoothed by a residual layer.
@@ III-F Loss Function
Since this paper focuses on exploring the potential of the Mamba architecture for CD tasks, we have optimized networks using commonly used loss functions in the field. Some tricks for improving accuracy like focal loss [62], deep supervision [23], multi-scale training/testing [6], etc. are not adopted. Nevertheless, we can find that the proposed architecture achieves very competitive and even SOTA performance on all three subtasks in the experimental section.
@@ III-F1 BCD
Since BCD can be regarded as a special semantic segmentation task [44], we directly optimize the network using the cross-entropy loss as
@@ III-F3 BDA
BDA contains a building location task and a damage classification task. The cross-entropy loss for each task can be expressed as follows
@@ IV Experiments And Discussions
| Dataset | Study site | Number of image pairs | Size | Resolution | Evaluation task |
| SYSU | Hong Kong | 20,000 | 256 × 256 | 0.5m | BCD (category-agnostic) |
| LEVIR-CD+ | Texas of the US | 985 | 1,024 × 1,024 | 0.5m | BCD (single-category) |
| WHU-CD | Christchurch, New Zealand | 1 | 32,508 × 15,354 | 0.3m | BCD (single-category) |
| xBD | 15 countries | 22,068 | 1,024 × 1,024 | <0.8m | BDA |
@@ IV-A1 SYSU-CD [19]
This dataset is category-agnostic CD dataset, which introduces a comprehensive collection of 20,000 pairs of 0.5 meter/pixel resolution aerial images from Hong Kong, spanning 2007 to 2014, to advance the field of CD. This dataset is distinguished by its focus on urban and coastal changes, featuring high-rise buildings and infrastructure developments, areas where CD poses significant challenges due to shadow and deviation effects. For effective deep learning (DL) application, the dataset is structured into training, validation, and test sets following a 6:2:2 ratio, with 20,000 patches of 256 × 256. Diverse change scenarios such as urban construction, suburban expansion, groundwork, vegetation changes, road expansion, and sea construction, are covered in this dataset.
@@ IV-A2 WHU-CD [64]
The WHU-CD dataset, a subset of the larger WHU Building dataset, is tailored for the building CD task. It comprises two aerial datasets from Christchurch, New Zealand, captured in April 2012 and 2016, with a spatial resolution of 0.3 meter/pixel. This dataset is particularly focused on detecting changes in large and sparse building structures. Aerial image capatured in 2012 cover an area of 20.5 km², featuring 12,796 buildings, while the image captured in 2016 shows an increase to 16,077 buildings within the same area, reflecting significant urban development over the four-year period. We follow the official protocol to split the dataset into training (21,243 15,354) area and testing (11,265 15,354) area for our experiments.
@@ IV-A3 LEVIR-CD+ [65]
This advanced version of the LEVIR-CD dataset is a building CD dataset. It comprises 985 pairs of very high-resolution images at 0.5 meter/pixel, each with dimensions of 1024 1024 pixels. Spanning a time interval of 5 to 14 years, these multi-temporal images document significant building construction changes. It also encompasses a wide array of building types, including villa residences, tall apartments, small garages, and large warehouses, with a focus on both the emergence of new buildings and the decline of existing structures. The final dataset features 31,333 instances of building changes, making LEVIR-CD+ a valuable benchmark for evaluating CD methodologies.
@@ IV-A4 SECOND [20]
The SECOND dataset introduces a collection of 4,662 pairs of 512 512 aerial images at 0.5-3 meter/pixel resolution, annotated for SCD in cities such as Hangzhou, Chengdu, and Shanghai. It focuses on six primary land cover classes: non-vegetated ground surfaces, trees, low vegetation, water, buildings, and playgrounds to capture a wide range of geographical changes. Annotations are done by experts, ensuring high accuracy at the pixel level. The dataset uses land-cover map pairs and non-change masks to accurately represent change, facilitating the differentiation of changes from unchanged pixels within the same class. This methodological approach ensures its utility in addressing SCD challenges by providing a diverse and accurately labeled dataset for model training and evaluation.
@@
The xBD dataset, pivotal for advancing BDA in disaster response, integrates pre- and post-disaster satellite imagery with over 850,000 building annotations across 45,362 . It stands as the largest dataset of its kind, enriched by environmental data and detailed damage levels for a broad spectrum of disaster events. The dataset adheres to stringent criteria, featuring high-resolution imagery (better than 0.8 meter/pixel), encapsulating multiple levels of damage (no damage, minor damage, major damage and destroyed), and covering a diverse range of disaster types (earthquake, tsunami, storm, wildfire, volcano and flood). By including both damaged and undamaged sites, xBD is crafted to foster the creation of adaptive models that effectively navigate the complexities of various disaster scenarios.